# A review of regularised estimation methods and cross-validation in spatiotemporal statistics

### v1

@@ 5 Cross validation under spatiotemporal dependence
For stationary time series, cross-validation (CV) is applied by segmenting the time series into subsets. Two main strategies are commonly used: (1) $k$-fold blocked subsets CV, where each subset is the test set in turn, and the rest are for training and forecasting (Bergmeir and Benítez,, 2012; Bergmeir et al.,, 2014, 2018), and (2) last-block CV or forward validation CV, where only the final block is the test set, and previous blocks are for training (Hjorth and Hjort,, 1982). The key distinction between the two approaches lies in preserving temporal order. In the last-block strategy, the model is never tested on past data relative to the training data. For instance, an $hv$-block method can be used to maintain independence between training and test sets, excluding a window of $h$ observations before and after the test set (Racine,, 2000). Moreover, both strategies allow choosing among several combinations of forecast horizons and updating schemes for the two samples (Tashman,, 2000, e.g., fixed-origin, rolling-origin and recalibration) to compute forecasting accuracy metrics (Bergmeir and Benítez,, 2012). A graphical synthesis of the two strategies is reported in Figure 1. Further CV strategies for temporal data can be found in Jiang and Wang, (2017) (Markov-CV and partitioned-CV), Cerqueira et al., (2020) (prequential approach), and Cerqueira et al., (2017) (Monte Carlo replications last-block approach). In the case of non-stationary time series, the previous CV schemes can be strongly misleading, as the unknown future may differ from the training sample, the test sample, or both. In this case, Bergmeir and Benítez, (2012) suggest using a weighted overlapping approach in which the whole series is used in training and testing steps.
@@
Figure 1: Temporal CV schemes. Blue blocks represent points from the time series in the training set, and red blocks represent points in the test set. Black blocks are neighbourhood observations to the test set and omitted from the training set.
Spatial data share challenges regarding autocorrelation with time series, but spatial data involves at least two-dimensional coordinates, whereas time is one-dimensional (Roberts et al.,, 2017). To address these issues, spatial CV methods and spatial variable selection techniques, like recursive and forward spatial feature selection, can be employed (Meyer et al.,, 2018, 2019). In addition, Schratz et al., (2019) demonstrated that hyperparameter tuning in machine learning approaches is less sensitive to spatial structure, but they recommend spatial cross-validation to ensure unbiased predictive performance.
@@
Moreover, Meyer and Pebesma, (2021) introduced the “area of applicability” (AOA) concept, defining the model’s valid geographical area based on training data. That is, the maximum distance (without outliers further apart than 1.5 interquartile ranges) in the covariate space of the training data defines the AOA in the prediction space. All predictions which are further apart than the maximum distance are marked as outside the AOA. The AOA will also account for the geographical distance if geographical coordinates are included in the covariates. In this way, the idea prevents predicting new geographic spaces with conditions that are very different from the training data, where the models can dramatically fail (Meyer and Pebesma,, 2022).
@@
The time-series block CV structure can be easily extended to spatial data by partitioning observations into spatial blocks. These blocks can be created either by dividing the entire space into cells for gridded data or by establishing spatial buffers around the data collection points for network or point data (Roberts et al.,, 2017). In Figure 2, we represent examples of spatial buffering both using grids and point data. In the latter, when the CV is performed by iteratively eliminating one location (and the neighbours within the buffer) at a time, we refer to spatial leave-one-location-out (SLOO) CV (Gasch et al.,, 2015; Meyer et al.,, 2018, 2019). As per LeRest2014, SLOO yields a criterion similar to the AIC but without accounting for spatial autocorrelation, meaning it produces the same output as AIC-based model selection in this context. However, when spatially correlated variables are in the model, AIC may not select the right covariates, whereas SLOO performs better. Additionally, spatial blocking can be applied to point patterns by assigning each location to its corresponding training polygon.
Figure 2: Spatial buffering for point data (left panel) and for gridded data (right panel).
@@
Alternative approaches to spatial blocking have been proposed in recent geostatistics literature. For instance, geographical partitioning provides a valid alternative via $k$-means algorithm (Brenning,, 2012). Provided a fixed number of partitions $k$, the clustering algorithm partitions the spatial location into non-overlapping clusters based on geodesic distance. Unfortunately, the simple $k$-means algorithm does not permit controlling the number of points in each partition, leading to potential heterogeneous partitions. To overcome this problem, Wang et al., (2021) proposed a heuristically modified k-means algorithm favouring partitions with similar elements. Also, see the inverse sampling-intensity weighting system by de Bruin et al., (2022) in which one can assign more weight to observations in sparsely sampled areas and less weight to observations in densely sampled areas to correct for estimation bias.
@@
When spatio-temporal data are considered, block-based CV schemes are obtained as a combination of the previously cited spatial and temporal blocking strategies. Following Meyer et al., (2018), we refer to time-block partitioning as Leave-Time-Out (LTO), point-in-space partitioning as Leave-Location-Out (LLO), and space-time partitioning as Leave-Location-and-Time-Out (LLTO). The three strategies are also called target-oriented to contrast the classical random approach. Specifically, the LTO partitions the spatiotemporal observations along the time axis into blocks of time series common to all locations (i.e., it performs training by eliminating all time instants assigned to the test block for all spatial locations and iterates over the time blocks). LLO does the same, but iterates w.r.t. the spatial blocks and considers the whole time series of each location. LTLO iterates over both temporal and spatial blocks (i.e., at a given iteration, the algorithm eliminates all observations of a certain spatial block and, for all other blocks, eliminates a common temporal block). Note that LLO coincides with the spatial blocking presented above, whereas LTO coincides with the temporal $k$-fold block strategy. In Figure 3, we show a schematic example of the three CV schemes for spatio-temporal data.
@@
Figure 3: Spatio-temporal CV schemes. Blue blocks represent points used for training the model, while red blocks represent points in the test set. Black blocks are neighbourhood observations to the test set and omitted from the training set.
@@
A further extension of the LLO approach is the Nearest Neighbour Distance Matching (NNDM) LOO CV introduced by Milà et al., (2022). This variant compares the nearest neighbour distance distribution function between the test and training data in the CV process to the nearest neighbour distance distribution function between the target prediction and training points. In practice, this is an alternative method to b-LLO in which the neighbours to be excluded are defined not by distance from the point but by the mismatch between the two Nearest Neighbour Distance distributions.
@@
Spatial and spatio-temporal resampling methods (CV and bootstrap) are implemented in many statistical software, such as R (R Core Team,, 2021). Among the others, see the packages sperrorest (Brenning,, 2012) (which implements distance-based K-means spatial partitioning); blockCV (Valavi et al.,, 2019), which implements block partitions and buffering for spatial data, as well as providing geostatistical tools for measuring spatial autocorrelation ranges in candidate covariates for model training and simplifying the choice of block and buffer sizes. It also offers an interactive tool for visualising spatial blocks as a function of folds and block/buffer sizes; CAST (Meyer et al.,, 2022), which implements the NNDM CV scheme as well as performing spatial variable selection to select suitable predictor variables according to their contribution to the spatial model performance; and mlr3spatiotempcv (Schratz et al.,, 2021), a package that is part of the mlr3 ecosystem (Lang et al.,, 2019) for the unified implementation of a wide range of ML models with feature and model selection tools and model evaluation capabilities. Specifically, mlr3spatiotempcv implements $k$-fold temporal and spatial blocking partitioning with and without buffering described in Meyer et al., (2018), and resumes the other partitioning techniques used in sperrorest, blockCV, skmeans (Zhao and Karypis,, 2002, hierarchical agglomerative clustering algorithms by), and CAST.
6 Conclusion and directions for future research
@@
In the field of spatial statistics and spatial econometrics, this comprehensive review has demonstrated the vital role of penalised methods in coping with the growing complexity of modern spatial data. With the increasing availability of geo-referenced data in various formats and types, the demand for adaptable, interpretable, and efficient modelling approaches becomes increasingly evident. This is where regularisation techniques step in, emerging as versatile tools for model selection, dimensionality reduction, and exploring spatial dependencies for classic statistical models. The advantage of interpretability sets them apart from the often enigmatic nature of deep learning models.
@@
In our review paper, we have presented the landscape of different regularisation methods, from the nuances of shrinkage to the mechanisms of penalisation strategies. We have underlined their practicality and effectiveness in the statistical modelling of geospatial data. We also briefly looked at Bayesian regularised estimation methods. Since the regularised estimation methods require the choice of a penalty term, which is usually done by cross-validation, we have also summarised cross-validation methods that can be used in the case of spatiotemporal dependence.
@@
There is a need for a more extensive exploration of regularised methods in spatial econometrics, particularly in estimating weighting matrices. This matrix is usually assumed to be known, which is rarely the case in practice, but allows the results to be interpreted in a geographical sense. To enhance the interpretability of these estimated matrices, one avenue to consider is using traditional distance-based weighting matrices as shrinkage targets, thereby enabling a geographical interpretation. Generally, the quest for enhanced computational efficiency in the application of regularisation methods for large-scale geospatial applications remains a pertinent concern in both fields.

### v2

@@ 5.1 Cross-validation in time
Cross-validation is usually applied for stationary time series by segmenting the time series into contiguous subsets to preserve the temporal correlation structure characterizing the observations. Two main strategies are commonly used: (1) $k$-fold blocked subsets CV, where each subset is treated as the test set in turn, and the rest are used for training and forecasting (Bergmeir and Benítez,, 2012; Bergmeir et al.,, 2014, 2018); and (2) last-block CV or forward validation CV, where only the final block is the test set, and previous blocks are used for training (Hjorth and Hjort,, 1982). The key distinction between the two approaches lies in preserving temporal order. Indeed, while the former preserves the natural order of temporal observations by forecasting future values using past observations (recall that the model is never tested on past data relative to the training data), the latter uses both past and future values to predict the current test set. In both cases, an $hv$-block method can be used to maintain independence between training and test sets, excluding a window of $h$ observations before and after the test set (Racine,, 2000). Moreover, both strategies allow choosing among several combinations of forecast horizons and updating schemes for the two samples (Tashman,, 2000, e.g., fixed-origin, rolling-origin and recalibration) to compute forecasting accuracy metrics (Bergmeir and Benítez,, 2012). A graphical synthesis of the two strategies is reported in Figure 3. Further CV strategies for temporal data can be found in Jiang and Wang, (2017) (Markov-CV and partitioned-CV), Cerqueira et al., (2020) (prequential approach), and Cerqueira et al., (2017) (Monte Carlo replications last-block approach).
Moreover, time series data are often characterised by non-stationarity of several typologies (Hewamalage et al.,, 2023). Sources of non-stationarity are seasonality, trends (both deterministic and stochastic), structural breaks or heteroskedasticity. Also, data can be far from Gaussianity due to heavy tails or outliers. In the case of non-stationary time series, the previous CV schemes can be strongly misleading, as the unknown future may differ from the training sample, the test sample, or both. Potential meaningful data splitting strategies include the use of a weighted overlapping approach in which the whole series is used in training and testing steps (Bergmeir and Benítez,, 2012) or the use of out-of-sample repeated holdout procedures applied in multiple testing periods (Cerqueira et al.,, 2020). For an extensive discussion of the role of partitioning schemes for temporal data in a forecasting context, see the recent paper by Hewamalage et al., (2023) in which detailed guidelines on the most appropriate strategy to implement based on the problem to be addressed and the relevant characteristics of the data at hand are provided.
@@
Figure 3: Temporal CV schemes. For a given iteration fold, blue blocks represent time points in the training set, and red blocks represent time points in the test set. Black blocks are the buffer points omitted from both the training and test sets.
5.2 Cross-validation in space
@@
While spatial data share similar challenges related to autocorrelation with time series, the former involves at least two-dimensional coordinates, whereas time is one-dimensional (Roberts et al.,, 2017). In addition, Schratz et al., (2019) demonstrated that hyperparameter tuning in machine learning approaches is less sensitive to spatial structure, but they recommend spatial cross-validation to ensure unbiased predictive performance. Moreover, Meyer and Pebesma, (2021) introduced the “area of applicability” (AOA) concept, defining the model’s valid geographical area based on training data. That is, the maximum distance (without outliers further apart than 1.5 interquartile ranges) in the covariate space of the training data defines the AOA in the prediction space. All predictions which are further apart than the maximum distance are marked as outside the AOA. The AOA also accounts for the geographical distance if geographical coordinates are included in the covariates. In this way, the idea prevents predicting new geographic spaces with conditions that are very different from the training data, where the models can dramatically fail (Meyer and Pebesma,, 2022).
@@
To address the above issues, spatial CV methods and spatial variable selection techniques, like recursive and forward spatial feature selection, can be employed (Meyer et al.,, 2018, 2019). The time-series block CV structure can be easily extended to spatial data by partitioning observations into spatial blocks. These blocks can be created either by dividing the entire space into cells for gridded data or by establishing spatial buffers between the training and test data (Roberts et al.,, 2017). In Figure 4, we represent examples of spatial buffering both using grids and point data. In the latter, when the CV is performed by iteratively eliminating one location (and the neighbours within the buffer) at a time, we refer to spatial leave-one-location-out (SLOO) CV (Gasch et al.,, 2015; Meyer et al.,, 2018, 2019). As per Le Rest et al., (2014), SLOO yields a criterion similar to the AIC but without accounting for spatial autocorrelation, meaning it produces the same output as AIC-based model selection in this context. However, when spatially correlated variables are in the model, AIC may not select the right covariates, whereas SLOO performs better. Additionally, spatial blocking can be applied to point patterns by assigning each location to its corresponding training polygon.
Figure 4: Spatial buffering for point data (left panel) and for gridded data (right panel). The red points represent the test locations, while the blue points represent training locations. Black points represent locations within the buffer of the test set, which are excluded from both the training and test.
@@
Alternative approaches to spatial blocking have been proposed in recent geostatistics literature. For instance, geographical partitioning provides a valid alternative via $k$-means algorithm (Brenning,, 2012). Provided a fixed number of partitions $k$, the clustering algorithm partitions the spatial locations into non-overlapping clusters based on geodesic distance. Unfortunately, the simple $k$-means algorithm does not permit controlling the number of points in each partition, leading to potential heterogeneous partitions. To overcome this problem, Wang et al., (2021) proposed a heuristically modified k-means algorithm favouring partitions with similar elements. One may also consider the inverse sampling-intensity weighting system proposed by de Bruin et al., (2022), in which one can assign more weight to observations in sparsely sampled areas and less weight to observations in densely sampled areas to correct for estimation bias. Even if not explicitly built for treating spatial and temporal CV tasks, potential extensions of the above-cited clustering approaches include the hierarchical spatiotemporal clustering with spatial constraints (Chavent et al.,, 2018) and spatially-clustered regression (Sugasawa and Murakami,, 2021).
5.3 Cross-validation in both space and time
@@
When spatiotemporal data are considered, block-based CV schemes are obtained as a combination of the previously cited spatial and temporal blocking strategies. Following Meyer et al., (2018), we refer to time-block partitioning as Leave-Time-Out (LTO), point-in-space partitioning as Leave-Location-Out (LLO), and space-time partitioning as Leave-Location-and-Time-Out (LLTO). The three strategies are also called target-oriented to contrast the classical random approach. Specifically, the LTO partitions the spatiotemporal observations along the time axis into blocks of time series common to all locations (i.e., it performs training by eliminating all time instants assigned to the test block for all spatial locations and iterates over the time blocks). LLO does the same but iterates with respect to the spatial blocks and considers the whole time series of each location. LLTO iterates over both temporal and spatial blocks (i.e., at a given iteration, the algorithm eliminates all observations of a certain spatial block and, for all other blocks, eliminates a common temporal block). Note that LLO coincides with the spatial blocking presented above, whereas LTO coincides with the temporal $k$-fold block strategy.
In Figure 5, we show a schematic example of the three above-mentioned target-oriented CV schemes for spatiotemporal data. Time stamps are reported on the rows for each panel, while spatial locations are on the columns. Regarding the latter, assume that each column represents a pair of longitude and latitude values and that stations are ordered according to some distance criterion. Cells marked in red represent values used in the test set, blue cells are the values used to train the model, and black cells are the buffering values used to separate training and test sets. According to the chosen scheme, the algorithm iterates over the rows (LTO), over the columns (LLO) or over the cells (LLTO). A geographical representation of the three algorithms is depicted in Figure 6, in which longitude and latitude are on the x-axis and y-axis, respectively, while each time stamp defines a specific panel.
Figure 5: Schematic matrix representation of the spatiotemporal target-oriented CV schemes. In the first row, the LLT algorithm iterates over the columns; in the second row, the LTO algorithm iterates over rows; in the third row, the LLTO algorithm iterates over spatiotemporal cells. Blue cells represent values used for training the model, while red cells represent points in the test set. Black blocks are observations lying within the spatiotemporal buffer that separate the test set from the training set.
@@
Figure 6: Schematic spatial representation of the spatiotemporal CV schemes. In LTO, the algorithm iterates over rows; in LLO, it iterates over columns; in LLTO, it iterates over cells. Blue cells represent values used for training the model, while red cells represent points in the test set. Black blocks are observations lying within the spatiotemporal buffer and separating the test set from the training set.
@@
A further extension of the LLO approach is the Nearest Neighbour Distance Matching (NNDM) LOO CV introduced by Milà et al., (2022). This variant compares the nearest neighbour distance distribution function between the test and training data in the CV process to the nearest neighbour distance distribution function between the target prediction and training points. In practice, this is an alternative method to b-LLO in which the neighbours to be excluded are defined not by distance from the point but by the mismatch between the two Nearest Neighbour Distance distributions. However, as the NNDM algorithm uses an LLO strategy, it is computationally intensive and cannot be used with large datasets. To overcome this issue, Linnenbrink et al., (2023) suggested a $k$-fold variant called the kNNDM algorithm.
5.4 Software for spatiotemporal cross-validation
Spatial and spatiotemporal CV methods are implemented in many statistical software packages, which we briefly review in this section. Here, we focus on the following main R packages (R Core Team,, 2023):
  1. 1.
    sperrorest (Brenning,, 2012) implements distance-based K-means spatial partitioning,
  2. 2.
    blockCV (Valavi et al.,, 2019) implements block partitions and buffering for spatial data, as well as providing geostatistical tools for measuring spatial autocorrelation ranges in candidate covariates for model training and simplifying the choice of block and buffer sizes. It also offers an interactive tool for visualising spatial blocks as a function of folds and block/buffer sizes,
  3. 3.
    CAST (Meyer et al.,, 2022, 2024) implements several spatiotemporal partitioning strategies including NNDM, kNNDMCV, LLO, LTO and LLTO schemes, as well as it allows performing spatial variable selection to select suitable predictor variables according to their contribution to the spatial model performance,
  4. 4.
@@ 5.4 Software for spatiotemporal cross-validation
    mlr3spatiotempcv (Schratz et al.,, 2021) is a package being part of the mlr3 ecosystem (Lang et al.,, 2019) which provides a unified implementation of a wide range of statistical learning models with feature and model selection tools and model evaluation capabilities. Specifically, mlr3spatiotempcv implements $k$-fold temporal and spatial blocking partitioning with and without buffering described in Meyer et al., (2018), and resumes the other partitioning techniques used in sperrorest, blockCV, skmeans (Zhao and Karypis,, 2002, hierarchical agglomerative clustering algorithms by), and CAST.
Moving to the Matlab environment, available software for spatiotemporal data partitioning includes the two-fold spatial CV strategy implemented in the DSTEM package (Wang et al.,, 2021) for the so-called hidden dynamic geostatistical model. Furthermore, spatiotemporal application of target-oriented CV schemes can be found in Otto et al., (2024) and Maranzano et al., (2023).
6 Summary and conclusion
@@
In the field of spatiotemporal statistics and spatial autoregression, this review has demonstrated the vital role of penalised methods in coping with the growing complexity of modern spatiotemporal data. With the increasing availability of geo-referenced data in various formats and types, the demand for adaptable, interpretable, and efficient modelling approaches becomes increasingly evident. This is where regularisation techniques step in, emerging as versatile tools for model selection, dimensionality reduction, and exploring spatial dependencies for classic statistical models. The advantage of interpretability sets them apart from the often enigmatic nature of deep learning models.
@@
In our review paper, we have presented a landscape of different regularisation methods, from the nuances of shrinkage to the mechanisms of penalisation strategies. We have underlined their practicality and effectiveness in the statistical modelling of geospatial data. We also briefly looked at Bayesian regularised estimation methods. Since the regularised estimation methods require the choice of a penalty term, which is usually done by cross-validation, we have also summarised cross-validation methods that can be used in the case of spatiotemporal dependence.
@@
There is a need for a more extensive exploration of regularised methods in spatial autoregression, particularly in estimating weighting matrices. This matrix is usually assumed to be known, which is rarely the case in practice, but allows the results to be interpreted in a geographical sense. To enhance the interpretability of these estimated matrices, one avenue to consider is using traditional distance-based weighting matrices as shrinkage targets, thereby enabling a geographical interpretation. Generally, the quest for enhanced computational efficiency in the application of regularisation methods for large-scale spatiotemporal applications remains a pertinent concern in both fields.